Filer Manufacturing has 9,404,869 shares of common stock outstanding. The current share price is $42.47, and the book value per share is $4.65. Filer Manufacturing also has two bond issues outstanding. The first bond issue has a face value of $49,730,426, has a 0.08 coupon, matures in 27 years and sells for 97 percent of par. The second issue has a face value of $68,173,795, has a 0.05 coupon, matures in 18 years, and sells for 97 percent of par. What is Filer's weight of equity on a market value basis? Enter the answer with 4 decimals (e.g. 0.2345)
Answer #1

Equity weight = 0.7774

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

в | C D E Market Value Nm 7 Capital Equity First Bond Second Bond $399,424,786.43 $48,238,513.22 $66,128,581.15 $513,791,880.

Cell reference -

Market Value Weight Capital Equity First Bond Second Bond =C8*C9 =C11* C12 =C14*C15 =SUM(C3:05) =C3/$C$6 =C4/$C$6 -C5/$C$6 Ou

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.